# A novel numerical framework for three-dimensional fully resolved simulation of freely falling particles of arbitrary shape

### v1

@@ Abstract
This article introduces a novel numerical framework designed to model the interplay between free-falling particles and their surrounding fluid in situations of high particle to fluid density ratio, typically exhibited by atmospheric particles. This method is designed to complement experimental studies in vertical wind tunnels to improve the understanding of the aerodynamic behavior of small atmospheric particles, such as the transport and sedimentation of volcanic particles, cloud ice crystals and other application areas. The solver is based on the lattice Boltzmann method and it addresses the numerical challenges, including the high density ratio and moderate to high Reynolds number, by using an immersed-boundary approach and a recursive-regularized collision model. A predictor-corrector scheme is applied for the robust time integration of the six-degrees-of-freedom (6DOF) rigid-body motion. Finally, the multi-scale nature arising from the long free-fall distances of a particle is addressed through a dynamic memory allocation scheme allowing for a virtually infinite falling distance. This tool allows for the simulation of particles of arbitrary shape represented by a triangularized surface. The framework is validated against the analytical and experimental data for falling spheres and ellipsoids, and is then applied to the case of an actual volcanic particle geometry, the shape of which is obtained from a 3D surface-contour scanning process. The physics of the free-fall of this particle is investigated and described, and its terminal velocity is compared against the experimental data measured with the 3D printed exemplars of the same particle.
@@
terminal velocity , free fall , air friction , lattice Boltzmann , immersed boundary method , Palabos , computational fluid dynamics , 6DOF
journal: Journal of Computational Science
@@ 1 Introduction
Particles moving in fluids are ubiquitous, from indoor air to the atmosphere and ocean to technical applications. With a few exceptions, such as small water droplets [1], such particles are almost universally non-spherical. Examples of such particles are volcanic ash [2], dust [3], pollen [4], ice crystals [5], marine snow [6] and microplastics [7], which mostly range in size from a few tens of micrometres to several centimetres (for the sake of brevity we exclude very small and large particles here). Their dynamics are influenced by their material properties such as density, size and shape, as well as by the properties of the ambient fluid and the flow characteristics. There is a long list of theoretical, experimental and numerical studies in the literature dealing with this topic [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. While experimental studies can provide the ground truth to guide and validate models, a reliable simulation can provide much more information about the particle dynamics that is not achievable in experiments. In recent years, several numerical methods have been proposed, which have demonstrated their capability in capturing some of the essential physical phenomena of such processes with high efficiency and accuracy [18, 19, 20, 21, 22].
The methods commonly rely on an immersed-particle philosophy, in which the fluid mesh is stationary, and mesh points are allocated both outside and inside the particle. The common implementations of this philosophy are the saturated-cell approach, which relies on a computation of the solid fraction across the interface of the moving particle [23, 24], or the immersed-boundary approach [19, 21], which simulates a fluid equation inside the solid as well. Both approaches avoid the need for costly remeshing of the fluid domain surrounding the moving particles. In this work, we focus on the immersed-boundary approach as one of the two common options in the Lattice Boltzmann modeling of rigid solid motion.
If the particle is considered rigid, its time evolution can be described by a model with six-degrees-of-freedom, or ‘6DOF’ for short, corresponding to three degrees of freedom for the position and three degrees of freedom for the orientation [25]. The resolution of the equations of motion relies on the computation of the total force and moment-of-force, which is obtained by integrating over an arbitrary particle shape [26].
A numerical framework is generally tailored for specific Reynolds numbers and particle to fluid density ratio regimes. Abundant literature exists for numerical simulation of particle suspensions at low Reynolds numbers and with density ratios close to 1 [27, 28, 29, 30]. These simulations have wide-ranging applications including sedimentation processes for particles immersed in fluids, where the particle density places them in a close-to-buoyant regime. To name another example, numerical simulation has made significant progress in representing blood components, including red blood cells or platelets, at the cellular level for biomedical process descriptions [31, 32, 33]. In such cases, the use of a solver capable of complete fluid-structure interaction is necessary due to the deformable nature of the blood cells.
@@
The particle dynamics is usually characterised by the Reynolds number of the particles $\mathrm{Re}_{T}=v_{\mathrm{T}}d/\nu$, which is defined as the relative velocity of the particles (here terminal velocity $v_{\mathrm{T}}$) multiplied by a characteristic length of the particles (here particle diameter $d$) and divided by the kinematic viscosity of the fluid $\nu$. In a turbulent flow, other dimensionless parameters also play a role. However, even in a laminar flow, it is not only the Reynolds number of the particles that is important, but also the ratio of the particle density to the fluid density. It has already been shown that for a Reynolds number $>$ 100, the terminal velocity increases with the density ratio [11, 15]. Recently, it was also found that the angular dynamics of particles with Reynolds number between 1-10 is also significantly influenced by the density ratio [17].
@@
One of the main difficulties in the experimental or numerical study of free-falling particles is their multi-scale nature. Length scales range from the size of the particles, e.g. a few hundred micrometres to centimetres, to drop distances that are ten or a thousand times the particle size to reach terminal velocity when a stable orientation is present or to collect sufficient statistics on the pseudo-terminal settling velocity of the particles when no stable orientation is present. Together with the need for high resolution to correctly resolve the shape features of the particles and the particle-fluid interactions, this problem can be fundamentally intractable given current computational power. A typical workaround calls for the limitation of the effective numerical domain size, but the close proximity to a no-slip wall representing the immutable reference frame (the ground) can in this case cause follow-up numerical and physical errors.
@@
We introduce in this work a novel numerical framework to model particle free-fall that addresses these issues. The framework is an extension of the Palabos library [34], a general-purpose computational fluid dynamics (CFD) code with a kernel based on the Lattice-Boltzmann method (LBM). The numerical stability of the flow is improved using the recursive-regularized collision model [35, 36]. This model also includes a subgrid-scale filtering capability achieved by projecting kinetic variables onto a Hermite-tensor subspace. This collision model therefore not only maintains stability but also ensures sufficient accuracy during local direct numerical simulation (DNS) to large-eddy simulation (LES) transition, which is expected to take place frequently in this flow regime. As demonstrated in this study, this LBM allows for the retrieval of fluid flow values with high precision to determine the force and moment of force acting on the surface of the immersed particle.
Owing to the adopted velocity discretization scheme, the LBM is traditionally limited to the use of structured meshes with uniform, cubic cells. Although this limitation is in practice overcome using mesh refinement on a hierarchic, octree-like mesh structure, such an extension is not used in the present work, which makes use of a fully uniform mesh. One of the consequences of the uniform mesh structure is that the mesh is not body-fitted for any object that has a non-rectangular shape or is moving.
@@
As a workaround, the LBM community has a well-established set of tools for accurately resolving boundary conditions for moving, immersed objects. Earlier methods, which were developed by Hu (1992) [37], Feng et al. (1994) [38, 39], Ladd (1994) [40, 41], Aidun (1995) [42], and Johnson (1995) [43], applied a staircased pattern to represent a particle boundary that followed the shape of the fluid solver’s mesh. They required a large particle size compared to the computational grid or mesh spacing to achieve stable and consistent results, or they needed to frequently remesh the fluid domain around the moving particles. Recently, a significant amount of research has been conducted to expand both conventional CFD approaches and the LBM in order to achieve a more precise representation of the surfaces of moving particles in fluids, without increasing the resolution of the computational domain [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. In separate research endeavors, the LBM has developed the ability to accurately represent curved boundaries over time [55].
In this paper, we concentrate on a multi-directional forcing technique that has was adapted for the LBM. In a research effort related to the present one, this approach has previously been used to model particles with density ratios close to unity [56]. The present work extends this framework to high density ratios and to particles of arbitrary shape. For this, the boundary representation is complemented with a time-integration scheme for the 6DOF of a rigid particle, leading to a complete computational framework for immersed particle motion, which is implemented in the LBM software library Palabos [34].
@@
The paper is divided into two parts. The first part outlines the computational framework for the particle free-fall experiment, which includes the recursive-regularized LBM for the fluid equations, the immersed boundary condition for the fluid-particle coupling, and a time-staggered predictor-corrector method for the time integration of the particle motion. To allow for a long free-fall distance without spending an excessive amount of memory, a dynamic memory allocation strategy is utilized. It involves a gradual removal of domains from the particle’s far wake and the addition of domains up front, permitting for the particle trajectory to be extended as often as necessary. The validity of the framework is confirmed in two simple cases, the time-dependent behavior of a sphere experiencing rotation-free free-fall and a displacement-free rotational motion. In the second part, the proposed numerical framework is used to investigate different scenarios with falling particles, such as spheroidal and irregular volcanic (lapilli-size) particles in a wide range of Reynolds number $\sim$ 5 - 10000 and a particle-to-fluid density ratio of $\sim$ 1000. The results are then compared with experimental measurements performed with the Göttingen Turret [17] and the UNIGE Vertical Wind Tunnel [57]. The numerical results show a high degree of agreement with the experimental observations, validating the numerical code.
2 Numerical method
2.1 Lattice Boltzmann method (LBM)
@@
Lattice Boltzmann models usually approximate the collision term by a perturbation around an equilibrium value, described by the Maxwell-Boltzmann distribution. This function depends on the macroscopic variables, such as, velocity and pressure, and is most often approximated by a polynomial of a given degree, which is constructed in a way to reproduce the selected velocity moments exactly. In this work, we select a fourth-order equilibrium term and model the collision with the recursive-regularized approach presented in [59], which is a more robust variation of the classical BGK algorithm (for Bhatnagar, Gross, and Krook, see [60]). The evolution rule reads
@@ 2.2 Immersed boundaries: direct forcing approach
An immersed boundary method (IBM) can represent the interaction of a fluid with an immersed solid, whose surface is not aligned with the Eulerian grid of the fluid. In particular, it can be used to simulate immersed moving solids, like the falling solid particles of this article, while the fluid mesh remains unchanged. The IBM gained some popularity through the approach developed by Peskin in [62] to study the heart valves. Since then, it has been developed further and applied to various areas. The basic idea of the IBM is that the surface of the solid object is approximated by a set of Lagrangian points, which are not required to coincide with the underlying Eulerian grid of the fluid. The fluid feels the presence of the solid by a singular force which is added to the momentum conservation equations. Regularized Dirac delta functions are used to transfer information from the Eulerian grid to the Lagrangian points, and vice versa. A very important variant of the classical IBM was proposed in [63]. This method, which is called the “direct forcing” approach, imposes a user-defined velocity at all points close to the immersed boundary. However, the no-slip boundary condition may not be effectively satisfied on each Lagrangian point due to conflicts with neighboring direct forcing terms. To solve this problem, Wang et al. [64] proposed the iterative “multi-direct forcing” approach, which can in principle be applied to any numerical method in CFD. The present work is based on the implementation of this technique in the LBM described in [65].
@@
For the next IBM cycle, we repeat from step $l=1$. In the present work, all solid surfaces are represented by triangularized meshes. The Lagrangian points ${\bm{X}}_{k}$ correspond to the vertices of this mesh, and the surface area $\delta{\bm{S}}_{k}$ corresponding to a vertex, is computed as one third of the sum of the areas of all triangles sharing this vertex.
@@ 2.3 Rigid particle motion
Table 1: Sequence of events for a time iteration of the coupled fluid – rigid-body algorithm. At each event, the second column specifies the time at which the input values are defined, and the subsequent columns the time to which the output values are taken. In the second column, the generic expression “$\mathrm{particle}$” is used to refer to the complete particle state, i.e. to the set of variables ${\bm{r}}$, ${\bm{d}}_{i}$, ${\bm{v}}$, and ${\bm{\omega}}$.
2.4 Coupling and time iteration
Figure 1: Simulation domain: three dimensional DNS domain for a particle falling in negative-y direction from the release position $h_{\mathrm{ini}}$ to the final position $h_{\mathrm{end}}$ is shown in the left. The domain height is $h$ in the y-direction and the lateral width is $w$, and a sponge zone $l_{\mathrm{sponge}}$ is allocated at the top portion of the domain. Once a particle reaches its y-direction limit $h_{\mathrm{end}}$, the red portion of the domain is copied together with the particle back to $h_{\mathrm{ini}}$, as shown in the right. See text for further explanations. The bottom boundary has the zero velocity $\bm{u}=0$ boundary condition, and the top boundary has constant fluid density boundary condition. The lateral boundaries have periodic boundary conditions.
@@ 2.5 Simulation setup
The setup and simulation principles of a numerical free-fall experiment is illustrated in Fig. 1. A particle of arbitrary shape is dropped either from rest or with an initial translational and / or angular velocity, and falls due to gravity in negative $y$-direction. The particle’s translational and rotational vectors are free and are impacted only by gravity and by air pressure and shearing forces. The top boundary, at the maximal $y$-coordinate, implements a constant fluid density, i.e. constant-pressure boundary condition. A sponge zone is additionally defined in the vicinity of the top boundary, to eliminate spurious pressure waves created in the domain by numerical artifacts. In this zone, the fluid viscosity is increased linearly to reach a relaxation time $\tau=1$ on the top boundary. The bottom boundary implements a no-slip condition. The lateral boundaries are periodic.
@@
It should be noted that for the largest simulated particles, with a diameter larger than 1 cm, the distance traversed by a particle until it reaches its terminal velocity is typically three orders of magnitude larger than the particle diameter. Therefore, to avoid the allocation of an impractically large domain, the particle is periodically reinjected at its initial position, as illustrated in Fig. 1. During this copy, the particle population of a full fluid volume directly impacted by the particle motion (indicated in red on the right part of the figure) is copied along with the particle. Also, when the particle drifts close to the lateral boundaries of the domain, a similar fluid volume around the particle is copied and repositioned in the center of the domain. This area is typically elongated above the particle, especially in high-Reynolds flows with a fully developed wake.
@@
It should be pointed out that this iterative reinjection scheme cannot be replaced by a simple periodicity condition between the top and the bottom wall, as such a setup would represent an open system with ever-increasing kinetic energy, under the effect of gravity. Another setup with lateral no-slip wall (instead of a no-slip bottom) would not be acceptable either, because it would represent a motion in a channel with a cross-channel velocity profile that is not representative of particle free-fall in an unbounded fluid.
@@
The simulation setup is caracterized by the domain height $h$ and width $w$, the positions $h_{\mathrm{ini}}$ and $h_{\mathrm{end}}$ for the initial and final position of a particle (before reinjection), the length of the sponge zone $l_{\mathrm{sponge}}$ and the number of lattices, i.e. $N$, spatially discretizes domain width $w$. The kinematic fluid viscosity, $\nu$, the density ratio between the particle and the air, $\kappa$, the volume-equivalent spherical diameter of the particle, $d$, and particle elongation, $el$, and flatness, $fl$, are further provided to complete the specifications of a numerical free-fall experiment. $el=b/a$ is defined as the ratio of the intermediate length $b$ of the particle to its largest length $a$, and $fl=c/b$ is defined as the ratio of the smallest length $c$ of the particle to the intermediate length $b$ [66]. Numerical length parameters are non-dimensionalized with respect to the volume-equivalent spherical diameter $d$, e.g. $w^{*}=w/d$. In all simulations, gravity has a value of 9.81 m2/s. The parameters of all numerical experiments of Section 3 are summarized in Table 2.

### v2

# Palabos Turret: A Particle-Resolved Numerical Framework for Settling Dynamics of Arbitrary-Shaped Particles

@@ Abstract
Particles transported in fluids are everywhere, occurring for example in indoor air, the atmosphere, the oceans, and engineering applications. In this study, a novel three-dimensional numerical framework – the Palabos Turret is presented, which allows fully resolved simulations of the settling dynamics of heavy particles with arbitrary shapes over a wide range of particle Reynolds numbers. The numerical solver is based on the lattice Boltzmann method utilizing immersed-boundary approach and a recursive-regularized collision model to fully resolve the particle-fluid interactions. A predictor-corrector scheme is applied for the robust time integration of the six-degrees-of-freedom (6DOF) rigid-body motion. Finally, the multi-scale nature arising from the long free-fall distances of a particle is addressed through a dynamic memory allocation scheme allowing for a virtually infinite falling distance. This solver allows for the simulation of particles of any arbitrary shape. The proposed framework is validated using the analytical and experimental data of freely-falling spheres, ellipsoids, and an irregular particle in a wide range of Reynolds numbers between $5\times 10^{-1}$ and $4\times 10^{4}$. For different Reynolds numbers and particle shapes considered, the Palabos Turret shows excellent agreement compared to theoretical and experimental values with a median relative deviation of ±$\pm 1.5\%$ and a maximum deviation of ±$\pm 5\%$. The Palabos Turret enables an in-depth analysis of the translational and rotational dynamics of particles with complex geometries.
@@
terminal velocity; free-fall; lattice Boltzmann; immersed boundary method; irregular geometry; rotational dynamics
@@ 1 Introduction
Solid particles moving in fluids are ubiquitous, from indoor air to the atmosphere and ocean to engineering applications. Such particles are almost universally non-spherical. Examples of such particles are volcanic ash [1], dust [2], pollen [3], ice crystals [4], marine snow [5] and microplastics [6], which mostly range in size from a few tens of micrometres to several centimetres (for the sake of brevity we exclude very small and large particles here). Their dynamics are influenced by their material properties such as density, size and shape, as well as by the properties of the ambient fluid and the flow characteristics. There is a long list of theoretical, experimental and numerical studies in the literature dealing with this topic [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. While experimental studies can provide the ground truth to guide and validate models, a reliable simulation can provide much more information about the particle dynamics that is hardly achievable in experiments. In recent years, several numerical methods have been proposed, which have demonstrated their capability in capturing some of the essential physical phenomena of such processes with high efficiency and accuracy [17, 18, 19, 20, 21].
@@
The particle dynamics is usually characterised by the Reynolds number of the particles $\mathrm{Re}=v_{\mathrm{T}}d/\nu$, which is defined as the relative velocity of the particles (here terminal velocity $v_{\mathrm{T}}$) multiplied by a characteristic length of the particles (here particle diameter $d$) and divided by the kinematic viscosity of the fluid $\nu$. In a turbulent flow, other dimensionless parameters also play a role. However, even in a laminar flow, it is not only the Reynolds number of the particles that is important, but also the ratio of the particle density to the fluid density. It has already been shown that for a $\mathrm{Re}>$ 100, the terminal velocity increases with the density ratio [10, 14]. Recently, it was also found that the angular dynamics of particles with Reynolds number between 1-10 is also significantly influenced by the density ratio [16]. If the particle is considered rigid, its time evolution can be described by a model with six-degrees-of-freedom, or ‘6DOF’ for short, corresponding to three degrees of freedom for the position and three degrees of freedom for the orientation [22]. The accuracy of the equations of motion relies on the computation of the total force and moment-of-force, which is obtained by integrating over an arbitrary particle shape [23].
A numerical framework is generally tailored for specific Reynolds numbers and particle to fluid density ratio regimes. Abundant literature exists for numerical simulation of particle suspensions at low Reynolds numbers and with density ratios close to 1 [24, 25, 26, 27]. These simulations have wide-ranging applications including sedimentation processes for particles immersed in fluids, where the particle density places them in a close-to-buoyant regime. To name another example, numerical simulation has made significant progress in representing blood components, including red blood cells or platelets, at the cellular level for biomedical process descriptions [28, 29, 30]. In such cases, the use of a solver capable of capturing fluid-structure interaction is necessary due to the deformable nature of the blood cells. The methods commonly rely on an immersed-particle philosophy, in which the fluid mesh is stationary, and the mesh points are allocated both outside and inside the particle. The common implementations of this philosophy are the saturated-cell approach, which relies on a computation of the solid fraction across the interface of the moving particle [31, 32]; or the immersed-boundary approach [18, 20], which simulates a fluid equation inside the solid particle as well. Both approaches avoid the need for costly remeshing of fluid domain surrounding the moving particles.
@@
One of the main difficulties in the experimental or numerical study of freely-falling particles is their multi-scale nature. For this problem, the length scales range from the size of the particle (e.g. a few hundred micrometres to centimetres) to the vertical drop distance which is between ten to thousand times the particle size to reach the terminal velocity, when the particle settles with a preferred orientation. When no preferred orientation is present, a large vertical drop distance is also necessary to collect sufficient statistics on the pseudo-terminal settling velocity of the particle. Together with the need for high resolution to correctly resolve the shape features of the particle and the particle-fluid interactions, this problem can be fundamentally intractable given current computational power. A common workaround involves restricting the effective numerical domain size, but in this case, the proximity to a no-slip wall representing the fixed reference frame (the ground) may lead to subsequent numerical and physical errors.
@@
In this work, we introduce a novel numerical framework to model particle free-fall that addresses these issues. Our framework is developed on the immersed-boundary approach in the lattice Boltzmann modeling of rigid solid motion. This framework, referred hereafter as the ‘Palabos Turret’, is an extension of the Palabos library [33] – a general-purpose computational fluid dynamics (CFD) code with a kernel based on the lattice Boltzmann method (LBM). The numerical stability of the flow is improved using the recursive-regularized collision model [34, 35]. This model also includes a subgrid-scale filtering capability, achieved by projecting kinetic variables onto a Hermite-tensor subspace. This collision model therefore not only maintains stability but also ensures sufficient accuracy during local direct numerical simulation (DNS) to large-eddy simulation (LES) transition, which is expected to take place frequently in this flow regime. It will be demonstrated that the Palabos Turret allows for the retrieval of fluid flow values with high precision to determine the force and moment of force acting on the surface of the immersed particle.
@@
Owing to the adopted velocity discretization scheme, the LBM is traditionally limited to the use of structured meshes with uniform, cubic cells. While this limitation is typically addressed by applying mesh refinement on a hierarchical, octree-like mesh structure, such an approach is not employed in the Palabos Turret, which instead uses a fully uniform mesh. One of the consequences of the uniform mesh structure is that the mesh is not body-fitted for any object that has a non-rectangular shape or is moving. As a workaround, the LBM community has a well-established a set of tools for accurately resolving boundary conditions for the moving, immersed objects. Earlier methods, which were developed by Hu et al. (1992) [36], Feng et al. (1994) [37, 38], Ladd (1994) [39, 40], Aidun & Lu (1995) [41], and Johnson & Tezduyar (1995) [42], applied a staircase pattern to represent a particle boundary that followed the shape of the fluid solver’s mesh. They required a large particle size compared to the computational grid or mesh spacing to achieve stable and consistent results, or they needed to frequently remesh the fluid domain around the moving particles. Recently, a significant amount of research has been conducted to expand both conventional CFD approaches and the LBM in order to achieve a more precise representation of the surfaces of moving particles in fluids, without increasing the resolution of the computational domain [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. In separate research endeavors, the LBM community has developed the ability to accurately represent curved boundaries over time [54]. In this work, we concentrate on a multi-directional forcing approach, adapted for the LBM. Such approach has previously been used to model particles with density ratios close to unity [55]. Our present work extends this approach to high density ratios and to particles of arbitrary shape. For this, the boundary representation is complemented with a time-integration scheme for the 6DOF of a rigid particle, leading to a complete computational framework for immersed particle motion.
@@
The rest of this paper is divided into three parts. Section 2 outlines the details of the Palabos Turret for the particle free-fall experiment, which includes the recursive-regularized LBM for the fluid equations, the immersed boundary condition for the fluid-particle coupling, and a time-staggered predictor-corrector method for the time integration of the particle motion. To allow for a long free-fall distance without spending an excessive amount of memory, a dynamic memory allocation strategy is utilized. In Section 3, after domain and mesh convergence tests, the validity of the Palabos Turret is firstly confirmed in two simple cases, the time-dependent behavior of a sphere experiencing rotation-free free-fall and a displacement-free rotational motion. The proposed Palabos Turret is then used to investigate different scenarios with freely-falling particles, such as spheroidal and irregular volcanic (lapilli-size) particles in a wide range of Reynolds number $\sim$ 0.5 - 40000 and a particle to fluid density ratio of $\sim$ 1000. The numerical results show a high degree of agreement with the analytical solutions and the experimental observations, validating the proposed Palabos Turret. Finally, a conclusion is drawn in Section 4.
2 The Palabos Turret
2.1 Lattice Boltzmann method
@@
Lattice Boltzmann models usually approximate the collision term by a perturbation around an equilibrium value, described by the Maxwell-Boltzmann distribution. This function depends on the macroscopic variables, such as, velocity and pressure, and is most often approximated by a polynomial of a given degree, which is constructed in a way to reproduce the selected velocity moments exactly. In this work, we select a fourth-order equilibrium term and model the collision with the recursive-regularized approach presented in [57], which is a more robust variation of the classical BGK algorithm (Bhatnagar, Gross & Krook (1954) [58]). The evolution rule reads
@@ 2.2 Immersed boundaries: direct forcing approach
An immersed boundary method (IBM) can represent the interaction of a fluid with an immersed solid, whose surface is not aligned with the Eulerian grid of the fluid. In particular, it can be used to simulate immersed moving solids, like the falling solid particles of this article, while the fluid mesh remains unchanged. The IBM gained some popularity through the approach developed by Peskin (1972) [60] to study the heart valves. Since then, it has been developed further and applied to various areas. The basic idea of the IBM is that the surface of the solid object is approximated by a set of Lagrangian points, which are not required to coincide with the underlying Eulerian grid of the fluid. The fluid feels the presence of the solid by a singular force which is added to the momentum conservation equations. Regularized Dirac delta functions are used to transfer information from the Eulerian grid to the Lagrangian points, and vice versa. A very important variant of the classical IBM was proposed in [61]. This method, which is called the ‘direct forcing’ approach, imposes a user-defined velocity at all points close to the immersed boundary. However, the no-slip boundary condition may not be effectively satisfied on each Lagrangian point due to conflicts with neighboring direct forcing terms. To solve this problem, Wang et al. (2008) [62] proposed the iterative ‘multi-direct forcing’ approach, which can in principle be applied to any numerical method in CFD. The present work is based on the implementation of this framework in the LBM described in [63].
@@
For the next IBM cycle, we repeat from step $l=1$. In the present work, all solid surfaces are represented by triangulated meshes. The Lagrangian points ${\bm{X}}_{k}$ correspond to the vertices of this mesh, and the surface area $\delta{\bm{S}}_{k}$ corresponding to a vertex, is computed as one third of the sum of the areas of all triangles sharing this vertex.
@@ 2.4 Coupling and time iteration
Table 1: Sequence of events for a time iteration of the coupled fluid – rigid-body algorithm. At each event, the second column specifies the time at which the input values are defined, and the subsequent columns the time to which the output values are taken. In the second column, the generic expression ‘$\mathrm{particle}$’ is used to refer to the complete particle state, i.e. to the set of variables ${\bm{r}}$, ${\bm{d}}_{i}$, ${\bm{v}}$, and ${\bm{\omega}}$.
Figure 1: Simulation domain: 3D computational domain for a particle (depicted as a black ellipsoid) falling in negative-$y$ direction from the release position $h_{\mathrm{ini}}$ to the final position $h_{\mathrm{end}}$ is shown in the left. The domain height is $h$ in the $y$-direction and the lateral width is $w$, and a sponge zone $l_{\mathrm{sponge}}$ is allocated at the top portion of the domain. Once a particle reaches its $y$-direction limit $h_{\mathrm{end}}$, the red portion of the domain is copied together with the particle back to $h_{\mathrm{ini}}$, as shown in the right. These values are specified in Table 2 or in corresponding results subsections for different cases considered in this study. See text for further explanations. No-slip boundary condition $\bm{u}=0$ and constant pressure are specified at the bottom and top boundaries, respectively. The lateral boundaries (i.e. all side walls parallel to $g$) have periodic boundary conditions.
@@ 2.5 Simulation setup
The setup and simulation principles of a numerical free-fall experiment is illustrated in Fig. 1. A particle of arbitrary shape is dropped either from rest or with an initial translational and / or angular velocity, and falls due to gravity in negative $y$-direction. The particle’s translational and rotational vectors are free and are impacted only by gravity $g$ and by air pressure and shearing forces. The top boundary, at the maximal $y$-coordinate, implements a constant fluid density, i.e. constant-pressure boundary condition. A sponge zone is additionally defined in the vicinity of the top boundary, to eliminate spurious pressure waves created in the domain by numerical artifacts. In this zone, the fluid viscosity is increased linearly to reach a relaxation time $\tau=1$ on the top boundary. The bottom boundary implements a no-slip condition. The lateral boundaries are periodic.
@@
It should be noted that the distance traversed by a particle until it reaches its terminal velocity is typically several orders of magnitude larger than the particle diameter. Therefore, to avoid the allocation of an impractically large domain, the particle is periodically re-injected at its initial position together with a copy of a portion of the fluid domain around it, as illustrated in red rectangle in Fig. 1. This copied domain contains the full fluid volume directly impacted by the particle motion, with the full width of $w$ and $\sim 60\%$ of the $h$. Also, when the particle drifts close to the lateral boundaries of the domain, a similar fluid volume around the particle is copied and re-positioned in the center of the domain. This area is typically elongated above the particle, especially in high-Reynolds flows with a fully developed wake.
@@
It should be pointed out that this iterative re-injection scheme cannot be replaced by a simple periodicity condition between the top and the bottom wall, as such a setup would represent an open system with ever-increasing kinetic energy, under the effect of gravity. Another setup with lateral no-slip wall (instead of a no-slip bottom) would not be acceptable either, because it would represent a motion in a channel with a cross-channel velocity profile that is not representative of particle free-fall in an unbounded fluid.
@@
The simulation setup is characterized by the domain height $h$ and width $w$, the positions $h_{\mathrm{ini}}$ and $h_{\mathrm{end}}$ for the initial and final position of a particle (before re-injection), the length of the sponge zone $l_{\mathrm{sponge}}$ and the number of lattices or the resolution $N$ spatially discretizes domain width $w$. To define the particle properties, the geometric parameters of the particle, diameter $d$ of a volume-equivalent sphere, particle density $\rho_{\mathrm{p}}$ and the initial conditions, such as initial orientation and velocity need to be provided. The fluid properties in the simulation are defined with the fluid density $\rho_{\mathrm{f}}$ and its kinematic viscosity $\nu$.